# Twist and higher modes of a complex scalar field at the threshold of collapse

### v1

@@ IV Numerical Results
A superposed oscillation (‘wiggle’), a familiar signature of DSS critical solutions, is clearly visible in this figure, with an estimated period of $\tilde{\Delta}\simeq 2.9\pm 0.2$. We determine this by separating the data into the two regions with distinct slopes, subtracting the respective linear trends, and fitting each segment with a sinusoidal function of the form $f(x)=A\sin(\omega x+b)+c$. The resulting periods are in reasonable agreement across both segments. Unlike in our previous work [2], where the DSS period of the phase space plot in Fig. 4 matched that of the central field values (see Fig. 3 and Table III in that paper), here the measured period in Fig. 8 is considerably longer than that in Fig. 6. Although small-scale features are already apparent by visual inspection, confirming whether the curve also contains a faster oscillation with period $\Delta\simeq 0.09$, as seen in the central values, would require a significantly more dense phase space sample of runs. Additionally, observe that due to the sparse sampling in parameter space of the $m=1$ case, Fig. 3 of the present work does not reveal any periodic wiggle at all. Therefore, we cannot conclude whether the difference in periods indicates a real discrepancy between the threshold descriptions.
@@ IV.4 Discussion on gravitational wave content and competition of thresholds
Figure 10: Log-log plot, from data of family III, of the ratio between the overall maxima of the Weyl invariant and the Kretschmann scalar, $W_{\mathrm{max}}/I_{\mathrm{max}}$, as a function of the distance to the threshold, $|a-a_{*}|$. The ratio remains approximately constant with some periodic features across the near-critical regime, with an average value $\ln(W_{\mathrm{max}}/I_{\mathrm{max}})\approx-0.71$.
@@ V Conclusions
In this work we performed the first axisymmetric simulations of massless complex scalar field collapse with twist and higher ($m=2$) angular modes. For this purpose we used the pseudospectral bamps code. The implementation of the $m$-cartoon symmetry reduction together with a generalization to our apparent horizon finder enabled us to probe the near-critical regime very closely for families with $m=1$ and $m=2$. For $m=1$ we recovered the DSS solution with $\Delta\simeq 0.42$ and $\gamma\simeq 0.11$, in agreement with [1]. For $m=2$ we identified for the first time a distinct DSS solution with smaller critical parameters, $\Delta\simeq 0.09$ and $\gamma\simeq 0.035$, showing that universality holds within fixed-$m$ families but that the critical exponents depend on the angular mode. We observe the presence of a periodic wiggle in the scaling plot of the $m=2$ case, with an estimated period $\tilde{\Delta}\simeq 2.9\pm 0.2$, notably larger than that of the central value plots. This represents an apparent difference with the threshold picture in [2], where both periods coincide. Given the limited sampling in parameter space, it is unclear whether this mismatch reflects a genuine physical difference or a resolution effect.
@@ Acknowledgments
We are grateful to T. Baumgarte, B. Brügmann, T. Giannakopoulos, C. Gundlach, D. Nitzschke, R. Pinto Santos and U. Sperhake for helpful discussions and feedback on various aspects of the work. We acknowledge financial support provided by FCT/Portugal through grants 2022.01324.PTDC, UID/99/2025. D. C. acknowledges support from the STFC Research Grant No. ST/V005669/1 “Probing Fundamental Physics with Gravitational-Wave Observations” and from the Leverhulme Trust Early Career Fellowship ECF-2025-424. H. R. R. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas-101052587. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or of the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A. V. V. thanks FCT for funding with DOI 10.54499/DL57/2016/CP1384/CT0090, as well as support from the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2022-134204-E, RED2022-134411-T, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). The figures in this article were produced with Plots.jl [48], Matplotlib [49], A. Khirnov’s nr_analysis_axi package [50] and ParaView [51, 52]. Several calculations were performed using the xAct package [53] for Mathematica. Numerical simulations were performed at the Leibniz Supercomputing Centre (LRZ), supported by the project pn36je. The authors thankfully acknowledge the computer resources, technical expertise and assistance provided by CENTRA/IST. Computations were also performed at the cluster “Baltasar-Sete-Sóis”.

### v2

@@ IV Numerical Results
A superposed oscillation (‘wiggle’), a familiar signature of DSS critical solutions, is clearly visible in this figure. We determine the DSS period, extracted from this oscillation, by separating the data into the two regions with distinct slopes and fitting each segment with a sinusoidal function of the form $f(x)=c+a\sin(\omega x+b)-2\gamma x$. We find frequencies $\omega_{1}\simeq 2.19\pm 0.16$ for the segment farer from the critical point and $\omega_{2}\simeq 2.18\pm 0.15$ for the near-critical one. These frequencies agree within the error bars derived from the fit. The resulting DSS periods are being calculated using the relation $\omega=\frac{\Delta}{2\gamma}$ taken from Eq. (18) in [9], since we expect that the maximum of the Ricci scalar scales according to the following relation
$\ln\left(R_{\mathrm{max}}\right)=c-2\gamma\ln\left(a_{*}-a\right)+f\left(2\gamma\ln\left(a_{*}-a\right)+b\right).$ (45)
In the first segment, this yields a DSS period
$\Delta_{1}=2\gamma_{1}\omega_{1}\simeq 0.096\pm 0.008,$ (46)
while in the near-critical segment, we obtain a DSS period
$\Delta_{2}=2\gamma_{2}\omega_{2}\simeq 0.15\pm 0.01.$ (47)
Surprisingly, the DSS segment of the segment farer from the threshold is in close agreement with the one measured by the central field values of the near-critical solution, $\Delta\simeq 0.09$. The DSS period of the near-critical segment on the other hand is almost double that value, owing to the almost doubled value of the relevant slope. Additionally, observe that due to the sparse sampling in parameter space of the $m=1$ case, Fig. 3 of the present work does not reveal any periodic wiggle at all. Overall, we believe that the threshold picture remains consistent with that of the Choptuik solution where universality and DSS manifest in the spacetime.
@@ IV.4 Discussion on gravitational wave content and competition of thresholds
Figure 10: Ratio between the overall maxima of the Weyl invariant and the Kretschmann scalar, $W_{\mathrm{max}}/I_{\mathrm{max}}$, from data of family III, as a function of the distance to the threshold, $|a-a_{*}|$. The ratio remains approximately constant with some periodic features across the near-critical regime, with an average value $\ln(W_{\mathrm{max}}/I_{\mathrm{max}})\approx-0.71$.
@@ V Conclusions
In this work we performed the first axisymmetric simulations of massless complex scalar field collapse with twist and higher ($m=2$) angular modes. We focused, as in earlier work, on initial data of pure $l=|m|$ harmonics in the scalarfield. For this purpose we used the pseudospectral bamps code. The implementation of the $m$-cartoon symmetry reduction together with a generalization to our apparent horizon finder enabled us to probe the near-critical regime very closely for families with $m=1$ and $m=2$. For $m=1$ we recovered the DSS solution with $\Delta\simeq 0.42$ and $\gamma\simeq 0.11$, in agreement with [1]. For $m=2$ we identified for the first time a distinct DSS solution with smaller critical parameters, $\Delta\simeq 0.09$ and $\gamma\simeq 0.035$, showing that universality holds within fixed-$m$ families but that the critical exponents depend on the angular mode. The phase space plot of the $m=2$ family exhibits two different slopes, but we only consider the one closer to the threshold for the calculation of the scaling exponent. In the same plot, we observe the presence of a periodic wiggle, with an estimated DSS period $\Delta_{1}\simeq 0.096\pm 0.008$ for the segment farer away from the threshold of collapse and a DSS period $\Delta_{2}\simeq 0.15\pm 0.01$ for the near-ciritical segment.
@@ Acknowledgments
We are grateful to T. Baumgarte, B. Brügmann, T. Giannakopoulos, C. Gundlach, D. Nitzschke, R. Pinto Santos and U. Sperhake for helpful discussions and feedback on various aspects of the work.
We acknowledge financial support provided by FCT/Portugal through grants 2022.01324.PTDC, UID/99/2025. D. C. acknowledges support from the STFC Research Grant No. ST/V005669/1, DiRAC projects ACTP284 and ACTP238, and from the Leverhulme Trust Early Career Fellowship ECF-2025-424. H. R. R. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas-101052587. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or of the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A. V. V. thanks FCT for funding with DOI 10.54499/DL57/2016/CP1384/CT0090, as well as support from the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2022-134204-E, RED2022-134411-T, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). The figures in this article were produced with Plots.jl [48], Matplotlib [49], A. Khirnov’s nr_analysis_axi package [50] and ParaView [51, 52]. Several calculations were performed using the xAct package [53] for Mathematica. Numerical simulations were performed at the Leibniz Supercomputing Centre (LRZ), supported by the project pn36je. The authors thankfully acknowledge the computer resources, technical expertise and assistance provided by CENTRA/IST. Computations were also performed at the cluster “Baltasar-Sete-Sóis”.